OpenStudy (anonymous):

x/-4 =-2/5 / this is a clearing fraction eqation can someone help me

jhonyy9 (jhonyy9):

x/-4 = -2/5 5x=8 x=8/5

OpenStudy (anonymous):

multiply everything by the LCM of 4 and 5, i.e. 20, and get \[-5x = -8; x = {8 \over 5}\]